Calculate the bond order for HF .
30 Mainly H H1s Energy F2p ALA IT 26 Mainly F 10 F2s
Answer #1

As shown in the diagram, H1s and F2s orbitals are far apart in energy and they do not interact. Hence the molecular orbital (MO), 1σ is nonbonding. The 1π orbitals does not have right symmetry to interact with H1s orbitals and they are nonbonding electrons on F atom. Non-bonding electrons does not contribute in the bond order. Here 2σ is the bonding orbital and 3σ is the antibonding orbital. Hence bond order

=1/2(bonding electrons - antibonding electrons) = 1/2(2-0) = 1